# Phase diagram of the one-dimensional three-state Potts model with an additional mean-field interaction

### v1

@@ Abstract
We derive the phase diagram of the one-dimensional three-state Potts model with an additional mean-field interaction in the canonical ensemble. The free energy is obtained by mapping the model onto the spin-$1$ Blume-Emery-Griffiths model and solving it by using an Hubbard-Stratonovich transformation combined with the transfer matrix method. A complex structure with lines of first-order transitions, two triple points and a critical point appears at finite temperature. The phase diagram is two-dimensional, since there are two adjustable parameters, the nearest-neighbour coupling $K$ and the temperature $T$. We show that the phase diagram does not present second-order phase transition lines, due to the fact that the order parameter is not a symmetry-breaking one. Quite remarkably, we are able to determine analytically one of the first-order phase-transition lines. We also prove that, when the nearest-neighbour coupling $K$ is large and negative, the first-order transition temperature becomes asymptotically independent of $K$.
@@ 1 Introduction
The interplay between short-range and long-range interactions in physical systems, such as those found in condensed matter and statistical mechanics, has long been a subject of research [1, 2, 3, 4, 5, 6, 7]. To formulate a minimal yet insightful framework for studying this interplay, one can use a discrete lattice spin model, for instance, a chain. This model would incorporate two types of interactions: nearest-neighbor couplings to represent short-range forces, and a long-range, mean-field-type component that connects each site to all others. A variant could be to have couplings that decay as a power law for large distances [7, 8] in presence of one or few additional short-range couplings. In the simplest setting with nearest-neighbour and all-to-all couplings, for classical Ising variables, the one-dimensional (1D) Hamiltonian of the model is typically written as
@@
where $K>0$ is the nearest-neighbour coupling constant, the sum runs over nearest-neighbor pairs $\langle i,j\rangle$, and $\delta_{\sigma_{i},\sigma_{j}}$ is the Kronecker delta, equal to $1$ if $\sigma_{i}=\sigma_{j}$ and 0 otherwise. It is clear that the case $q=2$ is equivalent to the Ising model, with just a redefinition of the coupling energies of a pair of spins that are in the same state or in different states. The Potts model exhibits a rich variety of behaviours depending on the number of states $q$ and the spatial dimensionality of the system. In two dimensions, the model undergoes a continuous phase transition for $q\leq 4$, while for $q>4$ the transition becomes first-order [12, 13]. Due to its versatility, the Potts model serves as a theoretical framework for studying not only magnetic systems, but also problems in percolation, lattice gauge theories, and combinatorics. In Ref. [12] the interested reader can find an exhaustive review dealing with these issues. We remind that in the purely mean-field Potts model, i.e., in the model in which the interaction between $\sigma_{i}$ and $\sigma_{j}$ is not restricted to nearest-neighbour spins, but it is the same for all pairs, the transition is of first-order for $q\geq 3$ [14].
@@
The $q$-state Potts model has a rich critical behaviour, either in the case of only nearest-neigbour coupling or only mean-field interaction. Moreover, it could show nontrivial effects induced by the competition between short-range and mean-field interactions as in the Nagle-Kardar model (1). In this paper we study a Potts-type extension of the Nagle-Kardar model. Such a system reveals novel phase transition phenomena, and provides further insight into how nonlocal interactions influence the collective behavior in multi-state systems, i.e., in systems where the single components, contrary to the case of Ising spins, can occupy more than two states. The crucial question is how a nearest-neighbour interaction affects the mean-field term when the number of spin states is larger than two. We focus on the case $q=3$, since already in this case the mean-field behaviour is different from that of the Ising model; in fact it is well known that the mean-field Ising model has a second-order phase transition. We consider both cases of positive (ferromagnetic) and negative (antiferromagnetic) nearest-neighbour couplings $K$. Notice that the Potts model with three states is different from the alternative generalization of the Nagle-Kardar model to the case in which $S$ is a spin $1$ variable (i.e., it assumes the values S=1,0,1S=1,0,-1). This latter model, when there are only all-to-all interactions, has a second-order phase transition (for a discussion of the spin-$1$ Nagle-Kardar model see Refs. [11, 15]).
@@
Summarizing our results, we underline that, in the presence of both the mean-field and the nearest-neighbour interaction the phase transitions remain first-order, and in the thermodynamic phase diagram (that, as we explain below, is two-dimensional) there are several lines of first-order transitions, with just a single critical point, that is of a very peculiar nature, being the end-point of three different first-order lines (we devote a section to its detailed study). Two of the first-order lines extend to infinite absolute values of the nearest-neighbour coupling constant, one to the positive side and the other to the negative side, yet with different characteristics. In the ferromagnetic case also the transition temperature increases without limits together with the coupling-constant (we are able to find an analytical expression for this transition line), while in the antiferromagnetic case the transition temperature remains finite, tending to an asymtpotic value that can be analytically obtained.
@@
The structure of the paper is as follows. In section 2 we explicitly give the Hamiltonian of the model and we explain how one can map it to the Hamiltonian of the spin-$1$ Blume-Emery-Griffiths model [16] in order to perform the analytical calculations. In this section we also anticipate that the equilibrium states never break completely the natural symmetry of the Potts model (in other words, the spontaneous symmetry breaking, always associated to the low temperature behaviour of the equilibrium states, is only partial), this property allowing to simplify the computations. In section 3 we show our results, presenting the phase diagram in the canonical ensemble and discussing several examples of the order parameter as a function of the temperature at different values of the nearest-neighbour coupling constant. Section 3.1 is devoted to the study of the phase diagram at large values of the coupling constant; on the other hand, section 3.2, gives a detailed analysis of the properties of the critical point. In section 4 we present our conclusions. Two appendices are related to a couple of side issues: in the first we give an a posteriori justification of the fact that the equilibrium states break only partially the symmetry of the model, while in the second we give an alternative evaluation of the first-order transition line for large negative values of the nearest-neighbour coupling constant.
@@ 2 The model
Performing one such transformation for each one of the two mean-field terms, the partition functions is obtained as:
@@
This expression shows that the free energy, and consequently the thermodynamic phase diagram, is obtained by a minimization problem with respect the two auxiliary variables $x$ and $y$. The values $x^{*}$ and $y^{*}$ realizing the minimum are, respectively, the values of the magnetization $m=(\sum_{i}S_{i})/N$ and of the quadrupole moment $q=(\sum_{i}S_{i}^{2})/N$ in the corresponding equilibrium state [17]. In the following, we will use $x$ and $y$ when referring to the auxiliary variables with respect to which the optimization has to be performed, and we will use $m$ and $q$, instead, when referring to the actual values of these variables, i.e, the magnetization and the quadrupole moment, in the equilibrium state. We have chosen to adhere to the usual convention to use the symbol $q$ for the quadrupole moment. In the literature this symbol is also usually employed to identify a $q$-state Potts model, with $q$ different spin states (and in fact in the Introduction we have used this symbol while referring to general Potts models), but this cannot generate confusion, since in this work we consider exclusively a three-state Potts model. We also point out that in the following we will refer to spin state as one of the states $a$, $b$ and $c$ (or 1-1, 0 and +1+1 in the BEG representation) in which the spin can be, while we will refer simply to state when talking about the equilibrium state of the whole system, as e.g., represented by the equilibrium values $m$ and $q$.
@@
The fact that the three spin states are characterized, as remarked above, only by being different spins states, without further quantitative property, has important consequences, that allow to simplify the optimization problem (17), arguing as follows. Let us denote as $(n_{a},n_{b},n_{c})$ the occupation fractions of the three spin states in a state of the system, being it in or out of equilibrium. Since there is a nearest-neighbour interaction, the three fractions $(n_{a},n_{b},n_{c})$ are not sufficient to fully characterize a state, and also the value of $\sum_{i}\delta_{S_{i}}\delta_{S_{i+1}}$ is necessary, but this is not relevant for our present argument. Because of the nature of the interactions in the Potts model, any state $(n_{a},n_{b},n_{c})$ is six-fold degenerate if the three fractions are all different; the other five states are obtained by the permutations of the fractions. On the other hand, the state will be three-fold degenerate if two of the fractions are equal, and it will be non-degenerate if the fractions are all equal to $1/3$. Restricting now to equilibrium states, the above identification, in the mapping to the BEG model, of the three spin states $a$, $b$ and $c$ of the Potts Hamiltonian with, respectively, the three spin states +1+1, 0 and 1-1 of the BEG Hamiltonian, implies that $m=n_{a}-n_{c}$ and $q=n_{a}+n_{c}$. It is also necessary to take into account the constraint $n_{a}+n_{b}+n_{c}=1$, so that actually only two of these three variables are necessary to specify the occupation fractions of the three spin states (coherently, in the BEG representation we need only the two parameters $m$ and $q$); assume that as these two variables we take the occupation fractions of the spin states $a$ and $b$, i.e., $n_{a}$ and $n_{b}$. Suppose now that an equilibrium state is characterized by the fractions $(n_{a}^{*},n_{b}^{*})$, with $n_{a}^{*}\neq n_{b}^{*}$ and also $n_{c}^{*}=1-n_{a}^{*}-n_{b}^{*}$ different from both $n_{a}^{*}$ and $n_{b}^{*}$. According to what said above concerning the degeneracy of the states, also the states $(n_{b}^{*},n_{a}^{*})$, $(n_{a}^{*},n_{c}^{*})$, $(n_{c}^{*},n_{a}^{*})$, $(n_{b}^{*},n_{c}^{*})$ and $(n_{c}^{*},n_{b}^{*})$ will be equilibrium states, i.e., the equilibrium state is six-fold degenerate. On the other hand, if $n_{a}^{*}\neq n_{b}^{*}=n_{c}^{*}$, then the equilibrium state will be only three-fold degenerate, the other two states being given by $(n_{b}^{*},n_{a}^{*})$ and $(n_{b}^{*},n_{b}^{*})$ (analogous situations occur when $n_{b}^{*}\neq n_{a}^{*}=n_{c}^{*}$ and when $n_{c}^{*}\neq n_{a}^{*}=n_{b}^{*}$). Going now to the BEG representation, these cases will translate in the following for the values of the magnetization $m$ and the quadrupole moment $q$. When $n_{a}^{*}\neq n_{b}^{*}\neq n_{c}^{*}\neq n_{a}^{*}$ the six degenerate equilibrium states will have $[m,q]$ equal to $\pm(n_{a}^{*}-n_{c}^{*}),n_{a}^{*}+n_{c}^{*}]$, $\pm(n_{a}^{*}-n_{b}^{*}),n_{a}^{*}+n_{b}^{*}]$ and $\pm(n_{b}^{*}-n_{c}^{*}),n_{b}^{*}+n_{c}^{*}]$. When $n_{a}^{*}\neq n_{b}^{*}=n_{c}^{*}$, the three degenerate equilibrium states will have $[m,q]$ equal to $\pm(n_{a}^{*}-n_{b}^{*}),n_{a}^{*}+n_{b}^{*}]$ and $0,2n_{b}^{*}]$ (and analogous cases when the two equal fraction are $n_{a}^{*}$ and $n_{b}^{*}$ or $n_{a}^{*}$ and $n_{c}^{*}$). Finally, when $n_{a}^{*}=n_{b}^{*}=n_{c}^{*}=1/3$, the non-degenerate equilibrium state will have $[m,q]=[0,2/3]$.
@@
The important point is the following. We have found that for all different values of $\beta$ and $K$ that we have tried, without exceptions, we fall in the second situation. In other words, the equilibrium state of the three-state Potts model is always one with at least two equal occupation fractions for the three spin states, so that the symmetry is broken only partially. This was already found for the purely mean-field ($K=0$) three-states Potts model [14]. This property allows to study the optimization problem (17) putting from the start $x$ equal to 0, then solving an optimization problem only with respect to $y$, simplifying considerably the numerical effort. It is not difficult to derive that, having found an equilibrium state with $[0,q]$, the other degenerate state must have $m_{1},q_{1}]=[\pm(3q/2-1),1-q/2]$. In Appendix A we provide an argument that shows the reason why the equilibrium states of this Potts model are all with at least two equal occupation fractions.
@@ 3 Results
As remarked above, the thermodynamic phase diagram is two-dimensional, with coordinates $(K,T)$, with the temperature $T=1/\beta$ (we choose units in which the Boltzmann constant $k_{B}$ is equal to $1$). For any given $K$ and $T$ the equilibrium value of the order parameter $y$ is obtained from the optimization problem (2). A necessary condition for a value $y$ to be the equilibrium one is the vanishing, for that value, of the derivative of $\widetilde{\phi}(\beta,K,y)$ with respect to $y$; however, this is not sufficient, since a vanishing derivative could be associated not only to a minimum, but also to a maximum and to an inflexion point. Moreover, there could be more than one local minimum, and the equilibrium would correspond to the absolute minimum. A peculiarity of our system is related to this issue. In fact, performing the partial $y$-derivative of $\widetilde{\phi}$, one finds that, regardless of the values of $\beta$ and $K$, at $y=2/3$ this derivative always vanishes. However, we also find that this value of $y$ corresponds to the global minimum only for sufficiently high temperatures $T$; in other words, for any given $K$, for all temperatures larger than a value dependent on $K$, the equilibrium value of the quadrupole moment is equal to $2/3$. This is consistent with the fact that at high temperatures we expect the three spin states to be equally populated, but this result shows that this is not achieved asymptotically with increasing temperature, but already at finite temperature. As our plots will show, furthermore, the change to an equilibrium value of $q$ equal to $2/3$ occurs through a first order phase transition.
Before giving the overall results for the phase diagram, we briefly consider what we expect for $T=0$. To this aim, it is convenient to consider the original Hamiltonian (3) and determine the configuration with the lowest energy among the configurations with equal number of spins in the states +1+1 and 1-1, since we have found that we can restrict ourselves to the configurations with $m=0$. For positive $K$ it is clear that the ground state is the one with $q=0$, i.e., with all the spins in the state 0. According to the Hamiltonian (3) its energy per spin is equal to (1/2)K-(1/2)-K. For negative $K$, on the other hand, we have to compare this energy with the one that pertains to the configuration with $q=1$ and with alternating spins in +1+1 and 1-1; its energy per spin is equal to (1/4)-(1/4). This is smaller than (1/2)K-(1/2)-K when K<(1/4)K<-(1/4). In conclusion, at $T=0$ there is a first order phase transition at K=(1/4)K=-(1/4) between a state with $q=0$ and all spins in the state 0, and a state with $q=1$ with alternating spins in the states +1+1 and 1-1.
@@
In the left plot of figure 1 we show the complete phase diagram, presenting lines of first-order phase transitions, in which the equilibrium value of the quadrupole moment $q$ undergoes a jump; we show examples of such behavior below. In the region of the phase diagram which is above all the lines the quadrupole moment $q$ is equal to $2/3$. The lines meet in the region that in the plot is marked with a small box. That region is blown up and shown in details in the right plot of the figure. In the left plot we see a line starting at $T=0$, at the value of $K$ equal to (1/4)-(1/4), obtained above as the value of first-order transition at vanishing temperature. The asymptotic behaviours of the other lines, the one extending to $K\to$-$\infty$ and the one extending to $K\to+\infty$, respectively, are obtained in section 3.1. We first comment, however, on the behaviour of the phase transition lines in the small region shown in the right plot. In this plot we see that there is a short line, the one going from the point denoted by TP1 to the point denoted by MCP, which is not visible at the scale of the left plot. Therefore in that small region there are actually three points, the two just mentioned and the point denoted by TP2, in which phase transition lines meet. The coordinates of these three points are the following. TP1: (0.19525,0.36820)(-0.19525,0.36820); MCP: (0.19342,0.38490)(-0.19342,0.38490); TP2: (0.17681,0.41328)(-0.17681,0.41328). Since the transitions are first-order, the points TP1 and TP2 are triple points. At first glance one would say the same for the point MCP; however, as we will prove in section 3.2, this point is actually a very peculiar critical point. The fact that we find points that are triple points and a critical point in our phase diagram is in agreement with the general behaviour of long-range systems analyzed in Ref. [18]: we have a Hamiltonian with one parameter, $K$, so that the phase diagram is two-dimensional; first-order transitions are associated with codimension 0 singularities, in turn associated to lines in the phase diagram, while triple and critical points are associated to codimension $1$ singularities, in turn associated to isolated points in the phase diagram. Also the absence of lines of second-order transitions can be understood in this framework. In fact, second-order transitions would be associated to codimension 0 singularities in the case in which they depend on an order parameter, such as $m$, with respect to which the Hamiltonian is symmetric (e.g., it is invariant for $m\to-m$). But our phase transitions depend on the order parameter $q$, the quadrupole moment, with respect to which the Hamiltonian does not have the analogous invariance, as already emphasized. In this case, second-order transitions are associated to codimension $1$ singularities, i.e., to points in the 2D phase diagram.
@@
Figure 2: The quadrupole moment $q$ as a function of the temperature $T$ for a fixed value of K=0.5K=-0.5 (left plot) and K=0.22K=-0.22 (right plot), showing the first-order transition with the corresponding jump of $q$. As expected from the phase diagram in figure 1, for K=0.5K=-0.5 there is one phase transition, while there are two phase transitions for K=0.22K=-0.22. Also, the plots show clearly that for all temperatures larger than that at the phase transition (or that at the last phase transition if there is more than one), $q$ is constant equal to $\frac{2}{3}$. This occurs for all values of $K$. We finally note that for $T\to 0$ the quadrupole moment tends to $1$ for K=0.5K=-0.5, while it tends to 0 for K=0.22K=-0.22, coherently with the fact that the ground state has $q=1$ for $K<-\frac{1}{4}$, while it has $q=0$ for $K>-\frac{1}{4}$. For a better visualization the plots, as well the left plot in figure 3 and the plots in figure 4, have a bottom and a top axis corresponding to values of $q$ slightly smaller than 0 and slightly larger than $1$, although these are obviously unphysical values.

### v2

@@ Abstract
We derive the phase diagram of the one-dimensional three-state Potts model with an additional mean-field interaction in the canonical ensemble. The free energy is obtained by mapping the model onto the spin-$1$ Blume-Emery-Griffiths model and solving it by using a Hubbard-Stratonovich transformation combined with the transfer matrix method. A complex structure with lines of first-order transitions, two triple points and a critical point appears at finite temperature. The phase diagram is two-dimensional, since there are two adjustable parameters, the nearest-neighbour coupling $K$ and the temperature $T$. We show that the phase diagram does not present second-order phase transition lines, due to the fact that the order parameter is not a symmetry-breaking one. Quite remarkably, we are able to determine analytically one of the first-order phase transition lines. We also prove that, when the nearest-neighbour coupling $K$ is large and negative, the first-order transition temperature becomes asymptotically independent of $K$.
@@ 1 Introduction
The interplay between short-range and long-range interactions in physical systems, such as those found in condensed matter and statistical mechanics, has long been a subject of research [1, 2, 3, 4, 5, 6, 7]. To formulate a minimal yet insightful framework for studying this interplay, one can use a discrete lattice spin model, for instance, a chain. This model would incorporate two types of interactions: nearest-neighbour couplings to represent short-range forces, and a long-range, mean-field-type component that connects each site to all others. A variant could be to have couplings that decay as a power law for large distances [7, 8] in presence of one or few additional short-range couplings. In the simplest setting with nearest-neighbour and all-to-all couplings, for classical Ising variables, the one-dimensional (1D) Hamiltonian of the model is typically written as
@@
where $K>0$ is the nearest-neighbour coupling constant, the sum runs over nearest-neighbour pairs $\langle i,j\rangle$, and $\delta_{\sigma_{i},\sigma_{j}}$ is the Kronecker delta, equal to $1$ if $\sigma_{i}=\sigma_{j}$ and 0 otherwise. It is clear that the case $q=2$ is equivalent to the Ising model, with just a redefinition of the coupling energies of a pair of spins that are in the same state or in different states. The Potts model exhibits a rich variety of behaviours depending on the number of states $q$ and the spatial dimensionality of the system. In two dimensions, the model undergoes a continuous phase transition for $q\leq 4$, while for $q>4$ the transition becomes first-order [12, 13]. Due to its versatility, the Potts model serves as a theoretical framework for studying not only magnetic systems, but also problems in percolation, lattice gauge theories, and combinatorics. In Ref. [12] the interested reader can find an exhaustive review dealing with these issues. We remind that in the purely mean-field Potts model, i.e., in the model in which the interaction between $\sigma_{i}$ and $\sigma_{j}$ is not restricted to nearest-neighbour spins, but it is the same for all pairs, the transition is of first-order for $q\geq 3$ [12, 14].
@@
The $q$-state Potts model has a rich critical behaviour, either in the case of only nearest-neighbour coupling or only mean-field interaction. Moreover, it could show nontrivial effects induced by the competition between short-range and mean-field interactions as in the Nagle-Kardar model (1). In this paper we study a Potts-type extension of the Nagle-Kardar model. Such a system reveals novel phase transition phenomena, and provides further insight into how nonlocal interactions influence the collective behaviour in multi-state systems, i.e., in systems where the single components, contrary to the case of Ising spins, can occupy more than two states. The crucial question is how a nearest-neighbour interaction affects the mean-field term when the number of spin states is larger than two. We focus on the case $q=3$, since already in this case the mean-field behaviour is different from that of the Ising model; in fact it is well known that the mean-field Ising model has a second-order phase transition. We consider both cases of positive (ferromagnetic) and negative (antiferromagnetic) nearest-neighbour couplings $K$. Notice that the Potts model with three states is different from the alternative generalization of the Nagle-Kardar model to the case in which $S$ is a spin $1$ variable (i.e., it assumes the values S=1,0,1S=1,0,-1). This latter model, when there are only all-to-all interactions, has a second-order phase transition (for a discussion of the spin-$1$ Nagle-Kardar model see Refs. [11, 15]).
@@
Summarizing our results, we underline that, in the presence of both the mean-field and the nearest-neighbour interaction the phase transitions remain first-order, and in the thermodynamic phase diagram (that, as we explain below, is two-dimensional) there are several lines of first-order transitions, with just a single critical point, that is of a very peculiar nature, being the end-point of three different first-order lines (we devote a section to its detailed study). Two of the first-order lines extend to infinite absolute values of the nearest-neighbour coupling constant, one to the positive side and the other to the negative side, yet with different characteristics. In the ferromagnetic case also the transition temperature increases without limits together with the coupling-constant (we are able to find an analytical expression for this transition line), while in the antiferromagnetic case the transition temperature remains finite, tending to an asymptotic value that can be analytically obtained.
@@
The structure of the paper is as follows. In section 2 we explicitly give the Hamiltonian of the model and we explain how one can map it to the Hamiltonian of the spin-$1$ Blume-Emery-Griffiths model [16] in order to perform the analytical calculations. In this section we also anticipate that the equilibrium states never break completely the natural symmetry of the Potts model (in other words, the spontaneous symmetry breaking, always associated to the low temperature behaviour of the equilibrium states, is only partial), this property allowing to simplify the computations. In section 3 we show our results, presenting the phase diagram in the canonical ensemble and discussing several examples of the order parameter as a function of the temperature at different values of the nearest-neighbour coupling constant. Section 3.1 is devoted to the study of the phase diagram at large values of the coupling constant; on the other hand, section 3.2 gives a detailed analysis of the properties of the critical point. In section 4 we present our conclusions. Two appendices are related to a couple of side issues: in the first we give an a posteriori justification of the fact that the equilibrium states break only partially the symmetry of the model, while in the second we give an alternative evaluation of the first-order transition line for large negative values of the nearest-neighbour coupling constant.
@@ 2 The model
Performing one such transformation for each one of the two mean-field terms, the partition function is obtained as:
@@
This expression shows that the free energy, and consequently the thermodynamic phase diagram, is obtained by a minimization problem with respect to the two auxiliary variables $x$ and $y$. The values $x^{*}$ and $y^{*}$ realizing the minimum are, respectively, the values of the magnetization $m=(\sum_{i}S_{i})/N$ and of the quadrupole moment $q=(\sum_{i}S_{i}^{2})/N$ in the corresponding equilibrium state [17]. In the following, we will use $x$ and $y$ when referring to the auxiliary variables with respect to which the optimization has to be performed, and we will use $m$ and $q$, instead, when referring to the actual values of these variables, i.e., the magnetization and the quadrupole moment, in the equilibrium state. We have chosen to adhere to the usual convention to use the symbol $q$ for the quadrupole moment. In the literature this symbol is also usually employed to identify a $q$-state Potts model, with $q$ different spin states (and in fact in the Introduction we have used this symbol while referring to general Potts models), but this cannot generate confusion, since in this work we consider exclusively a three-state Potts model. We also point out that in the following we will refer to spin state as one of the states $a$, $b$ and $c$ (or 1-1, 0 and +1+1 in the BEG representation) in which the spin can be, while we will refer simply to state when talking about the equilibrium state of the whole system, as e.g., represented by the equilibrium values $m$ and $q$.
@@
The fact that the three spin states are characterized, as remarked above, only by being different spins states, without further quantitative property, has important consequences, that allow to simplify the optimization problem (17), arguing as follows. Let us denote as $(n_{a},n_{b},n_{c})$ the occupation fractions of the three spin states in a state of the system, being it in or out of equilibrium. Since there is a nearest-neighbour interaction, the three fractions $(n_{a},n_{b},n_{c})$ are not sufficient to fully characterize a state, and also the value of $\sum_{i}\delta_{S_{i},S_{i+1}}$ is necessary, but this is not relevant for our present argument. Because of the nature of the interactions in the Potts model, any state $(n_{a},n_{b},n_{c})$ is six-fold degenerate if the three fractions are all different; the other five states are obtained by the permutations of the fractions. On the other hand, the state will be three-fold degenerate if two of the fractions are equal, and it will be non-degenerate if the fractions are all equal to $1/3$. Restricting now to equilibrium states, the above identification, in the mapping to the BEG model, of the three spin states $a$, $b$ and $c$ of the Potts Hamiltonian with, respectively, the three spin states +1+1, 0 and 1-1 of the BEG Hamiltonian, implies that $m=n_{a}-n_{c}$ and $q=n_{a}+n_{c}$. It is also necessary to take into account the constraint $n_{a}+n_{b}+n_{c}=1$, so that actually only two of these three variables are necessary to specify the occupation fractions of the three spin states (coherently, in the BEG representation we need only the two parameters $m$ and $q$); assume that as these two variables we take the occupation fractions of the spin states $a$ and $b$, i.e., $n_{a}$ and $n_{b}$. Suppose now that an equilibrium state is characterized by the fractions $(n_{a}^{*},n_{b}^{*})$, with $n_{a}^{*}\neq n_{b}^{*}$ and also $n_{c}^{*}=1-n_{a}^{*}-n_{b}^{*}$ different from both $n_{a}^{*}$ and $n_{b}^{*}$. According to what said above concerning the degeneracy of the states, also the states $(n_{b}^{*},n_{a}^{*})$, $(n_{a}^{*},n_{c}^{*})$, $(n_{c}^{*},n_{a}^{*})$, $(n_{b}^{*},n_{c}^{*})$ and $(n_{c}^{*},n_{b}^{*})$ will be equilibrium states, i.e., the equilibrium state is six-fold degenerate. On the other hand, if $n_{a}^{*}\neq n_{b}^{*}=n_{c}^{*}$, then the equilibrium state will be only three-fold degenerate, the other two states being given by $(n_{b}^{*},n_{a}^{*})$ and $(n_{b}^{*},n_{b}^{*})$ (analogous situations occur when $n_{b}^{*}\neq n_{a}^{*}=n_{c}^{*}$ and when $n_{c}^{*}\neq n_{a}^{*}=n_{b}^{*}$). Going now to the BEG representation, these cases will translate in the following for the values of the magnetization $m$ and the quadrupole moment $q$. When $n_{a}^{*}\neq n_{b}^{*}\neq n_{c}^{*}\neq n_{a}^{*}$ the six degenerate equilibrium states will have $[m,q]$ equal to $\pm(n_{a}^{*}-n_{c}^{*}),n_{a}^{*}+n_{c}^{*}]$, $\pm(n_{a}^{*}-n_{b}^{*}),n_{a}^{*}+n_{b}^{*}]$ and $\pm(n_{b}^{*}-n_{c}^{*}),n_{b}^{*}+n_{c}^{*}]$. When $n_{a}^{*}\neq n_{b}^{*}=n_{c}^{*}$, the three degenerate equilibrium states will have $[m,q]$ equal to $\pm(n_{a}^{*}-n_{b}^{*}),n_{a}^{*}+n_{b}^{*}]$ and $0,2n_{b}^{*}]$ (and analogous cases when the two equal fraction are $n_{a}^{*}$ and $n_{b}^{*}$ or $n_{a}^{*}$ and $n_{c}^{*}$). Finally, when $n_{a}^{*}=n_{b}^{*}=n_{c}^{*}=1/3$, the non-degenerate equilibrium state will have $[m,q]=[0,2/3]$.
@@
The important point is the following. We have found that for all different values of $\beta$ and $K$ that we have tried, we fall, without exceptions, in the second situation. In other words, the equilibrium state of the three-state Potts model is always one with at least two equal occupation fractions for the three spin states, so that the symmetry is broken only partially. This was already found for the purely mean-field ($K=0$) three-states Potts model [14]. This property allows to study the optimization problem (17) putting from the start $x$ equal to 0, then solving an optimization problem only with respect to $y$, simplifying considerably the numerical effort. It is not difficult to derive that, having found an equilibrium state with $[0,q]$, the other degenerate state must have $m_{1},q_{1}]=[\pm(3q/2-1),1-q/2]$. In Appendix A we provide an argument that shows the reason why the equilibrium states of this Potts model are all with at least two equal occupation fractions.
@@ 3 Results
As remarked above, the thermodynamic phase diagram is two-dimensional, with coordinates $(K,T)$, with the temperature $T=1/\beta$ (we choose units in which the Boltzmann constant $k_{B}$ is equal to $1$). For any given $K$ and $T$ the equilibrium value of the order parameter $y$ is obtained from the optimization problem (2). A necessary condition for a value $y$ to be the equilibrium one is the vanishing, for that value, of the derivative of $\widetilde{\phi}(\beta,K,y)$ with respect to $y$; however, this is not sufficient, since a vanishing derivative could be associated not only to a minimum, but also to a maximum and to an inflexion point. Moreover, there could be more than one local minimum, and the equilibrium would correspond to the absolute minimum. A peculiarity of our system is related to this issue. In fact, performing the partial $y$-derivative of $\widetilde{\phi}$, one finds that, regardless of the values of $\beta$ and $K$, at $y=2/3$ this derivative always vanishes. However, we also find that this value of $y$ corresponds to the global minimum only for sufficiently high temperatures $T$; in other words, for any given $K$, for all temperatures larger than a value dependent on $K$, the equilibrium value of the quadrupole moment is equal to $2/3$. This is consistent with the fact that at high temperatures we expect the three spin states to be equally populated, but this result shows that this is not achieved asymptotically with increasing temperature, but already at finite temperature. As our plots will show, furthermore, the change to an equilibrium value of $q$ equal to $2/3$ occurs through a first-order phase transition.
Before giving the overall results for the phase diagram, we briefly consider what we expect for $T=0$. To this aim, it is convenient to consider the original Hamiltonian (3) and determine the configuration with the lowest energy among the configurations with equal number of spins in the states +1+1 and 1-1, since we have found that we can restrict ourselves to the configurations with $m=0$. For positive $K$ it is clear that the ground state is the one with $q=0$, i.e., with all the spins in the state 0. According to the Hamiltonian (3) its energy per spin is equal to (1/2)K-(1/2)-K. For negative $K$, on the other hand, we have to compare this energy with the one that pertains to the configuration with $q=1$ and with alternating spins in +1+1 and 1-1; its energy per spin is equal to (1/4)-(1/4). This is smaller than (1/2)K-(1/2)-K when K<(1/4)K<-(1/4). In conclusion, at $T=0$ there is a first-order phase transition at K=(1/4)K=-(1/4) between a state with $q=0$ and all spins in the state 0, and a state with $q=1$ with alternating spins in the states +1+1 and 1-1.
@@
In the left plot of figure 1 we show the complete phase diagram, presenting lines of first-order phase transitions, in which the equilibrium value of the quadrupole moment $q$ undergoes a jump; we show examples of such behaviour below. In the region of the phase diagram which is above all the lines the quadrupole moment $q$ is equal to $2/3$. The lines meet in the region that in the plot is marked with a small box. That region is blown up and shown in details in the right plot of the figure. In the left plot we see a line starting at $T=0$, at the value of $K$ equal to (1/4)-(1/4), obtained above as the value of first-order transition at vanishing temperature. The asymptotic behaviours of the other lines, the one extending to $K\to$-$\infty$ and the one extending to $K\to+\infty$, respectively, are obtained in section 3.1. We first comment, however, on the behaviour of the phase transition lines in the small region shown in the right plot. In this plot we see that there is a short line, the one going from the point denoted by TP1 to the point denoted by MCP, which is not visible at the scale of the left plot. Therefore in that small region there are actually three points, the two just mentioned and the point denoted by TP2, in which phase transition lines meet. The coordinates of these three points are the following. TP1: (0.19525,0.36820)(-0.19525,0.36820); MCP: (0.19342,0.38490)(-0.19342,0.38490); TP2: (0.17681,0.41328)(-0.17681,0.41328). Since the transitions are first-order, the points TP1 and TP2 are triple points. At first glance one would say the same for the point MCP; however, as we will prove in section 3.2, this point is actually a very peculiar critical point. The fact that we find points that are triple points and a critical point in our phase diagram is in agreement with the general behaviour of long-range systems analyzed in Ref. [18]: we have a Hamiltonian with one parameter, $K$, so that the phase diagram is two-dimensional; first-order transitions are associated with codimension 0 singularities, in turn associated to lines in the phase diagram, while triple and critical points are associated to codimension $1$ singularities, in turn associated to isolated points in the phase diagram. Also the absence of lines of second-order transitions can be understood in this framework. In fact, second-order transitions would be associated to codimension 0 singularities in the case in which they depend on an order parameter, such as $m$, with respect to which the Hamiltonian is symmetric (e.g., it is invariant for $m\to-m$). But our phase transitions depend on the order parameter $q$, the quadrupole moment, with respect to which the Hamiltonian does not have the analogous invariance, as already emphasized. In this case, second-order transitions are associated to codimension $1$ singularities, i.e., to points in the 2D phase diagram.
@@
Figure 2: The quadrupole moment $q$ as a function of the temperature $T$ for a fixed value of K=0.5K=-0.5 (left plot) and K=0.22K=-0.22 (right plot), showing the first-order transitions with the corresponding jump of $q$. As expected from the phase diagram in figure 1, for K=0.5K=-0.5 there is one phase transition, while there are two phase transitions for K=0.22K=-0.22. Also, the plots show clearly that for all temperatures larger than that at the phase transition (or that at the last phase transition if there is more than one), $q$ is constant equal to $\frac{2}{3}$. This occurs for all values of $K$. We finally note that for $T\to 0$ the quadrupole moment tends to $1$ for K=0.5K=-0.5, while it tends to 0 for K=0.22K=-0.22, coherently with the fact that the ground state has $q=1$ for $K<-\frac{1}{4}$, while it has $q=0$ for $K>-\frac{1}{4}$. For a better visualization the plots, as well the left plot in figure 3 and the plots in figure 4, have a bottom and a top axis corresponding to values of $q$ slightly smaller than 0 and slightly larger than $1$, although these are obviously unphysical values.